1.2x-10+13+1=5x+11
2.8x/3-3=2x/3+15
3.12x/5+5=29-3x/5
4.2x/5-2=12x/5+11
5.3x/7-4=24x/7+8

Found 2 solutions by sdmmadam@yahoo.com, ali.m:
Answer by sdmmadam@yahoo.com(530)

1.2x-10+13+1=5x+11
-10+13+1-11= 5x - 2x (taking 2x to the right and bringing 11 to the left)
(change side then change sign)
(-10-11)+(13+1) = 3x (by additive commutativity and associativity)
-21+14 =3x
-7 = 3x
(-7)/3 = x (dividing both the sides by 3)
Therefor x = -7/3
Verification: Put x=-7/3 on the LHS. You get (-14/3) +4 = (-14+12)/3 = -2/3
RHS = 5x+11 = 5X(-7/3) + 11 = (33-35)/3 = -2/3 LHS
Therefore x = -7/3 is right
Note: The advantage of commutativity is you can pluck a term from some position on a side and place it in some other position on the same side along with the sign of course.(transplantation!)
2.8x/3-3=2x/3+15
(8x/3- 2x/3) = 15+3 (transposing, that is changing side)
(8x-2x)/3 = 18
6x/3 = 18
2x =18
x=9
Answer: x = 9
Verification: for x=9 LHS = 21 =RHS
3.12x/5+5=29-3x/5
12x/5+3x/5 = 29 - 5
(12x+3x)/5 = 24
15x/5 = 24
3x=24
x = 24/3 = 8
x=8
Verification: LHS = (96/5+5) = 121/5
RHS = (29-24/5) = (145-24)/5 = 121/5 = LHS
Therefore x = 8 is correct

4.2x/5-2=12x/5+11
2x/5 -12x/5 = 11 +2
(2x-12x)/5 = 13
(-10x)/5 = 13
-2x = 13
x = 13/(-2) = -13/2
x = -13/2
Verification: LHS = 2x/5-2 = -13/5 -2 = (-13-10)/5 = -23/5
RHS= 12x/5 + 11 = -78/5 +11 = (55-78)/5 = -23/5 = LHS
Therefore x = -13/2 is correct


5.3x/7-4=24x/7+8
3x/7 - 24x/7 = 8 +4
(3x-24x)/7 = 12
-21x/7 = 12
-(3X7x)/7 = 12
-3x = 12 (cancelling 7 in the nr and in the dr)
x = 12/(-3) = -4
Answer: x = -4
Verification: LHs = 3x/7-4 (3X(-4)/7 -4 = -12/7 -4 = (-28-12)/7 = -40/7
RHS = 24x/7+8 = 24X(-4)/7 + 8 = -96/7 + 8 = (56-96)/7 = -40/7 = LHS
Therefore x = -4

Answer by ali.m(1)

2x-10+13+1=5x+11